# Integrating large language models and active inference to understand eye movements in reading and dyslexia

### v2

@@ 1.2 Brief introduction to Active Inference
Active Inference is a framework that models an agent’s action-perception loop by minimizing variational free energy (Parr et al., 2022). The key objective is to minimize free energy, and to achieve this, the agent possesses a generative model (like the one illustrated in Figure 1), which captures the joint probability of the stochastic variables (hidden states and observations) using the formalism of probabilistic graphical models (Bishop and Nasrabadi, 2006). The agent’s perception involves making inferences about the hidden states based on the observed sensory inputs, while action involves selecting actions that can change the hidden states and optimize the model’s predictions. The generative model for active inference is defined as follows:
$P\left(o_{0:K},s_{0:K},u_{1:K},\gamma|\boldsymbol{\Theta}\right)=$
$P\left(\gamma|\boldsymbol{\Theta}\right)P\left(\pi|\gamma,\boldsymbol{\Theta}\right)P\left(s_{0}|\boldsymbol{\Theta}\right)\prod_{t=0}^{K}P\left(o_{t}|s_{t},\boldsymbol{\Theta}\right)P\left(s_{t+1}|s_{t},\pi_{t},\boldsymbol{\Theta}\right)$ (1)
@@
  • 2.
    The (transition) matrix $\mathbf{B}$ defines how hidden states evolve over time $t$, as a function of a control state (action) $u_{t}$; note that a sequence of control states $u_{1}$,$u_{2}$$\dots$,$u_{t}$$\dots$ defines an action policy (or policy for short) $\pi_{t}$.
  • 3.
    The matrix $\mathbf{C}$ is an a-priori probability distribution over observations (and typically encodes the agent’s preferences).
  • 4.
@@
In active inference, the process of perception involves estimating hidden states based on observations and previous hidden states. At the start of the simulation, the model has access to an initial state estimate $s_{0}$ through $\mathbf{D}$ and receives an observation $o_{0}$ that helps refine the estimate by using the likelihood matrix $\mathbf{D}$. Subsequently, for each time step $t=1,\dots,K$, the model infers its current hidden state $s_{t}$ by considering the transitions determined by the control state $u_{t}$, as specified in $\mathbf{B}$. Active inference employs an approximate posterior over (past, present, and future) hidden states and parameters ($s_{0:K},u_{1:K},\gamma$). This approach utilizes a factorized form of variational inference, which corresponds to a framework developed in physics known as mean field theory (Parisi, 1988). Using a mean field approximation, namely assuming that all variables are independent, the factorized approximated posterior can be expressed as:
@@
where the sufficient statistics are encoded by the expectations $\boldsymbol{\mu}=\left(\mathbf{\tilde{s}}^{\boldsymbol{\pi}},\boldsymbol{\pi},\boldsymbol{\gamma}\right)$, with $\mathbf{\tilde{s}}^{\boldsymbol{\pi}}=\mathbf{\tilde{s}}_{0}^{\boldsymbol{\pi}},\dots,\mathbf{\tilde{s}}_{K}^{\boldsymbol{\pi}}$. Following a variational approach, the distribution in (2) best approximates the posterior when its sufficient statistics $\mu$ minimise the free energy of the generative model, see (Parr et al., 2022). This condition holds when the sufficient statistics are:
@@
Action selection in the active inference framework involves choosing a policy, represented by a sequence of control states $u_{1},u_{2},\dots,u_{t}$, that is expected to minimize free energy most effectively in the future. The policy distribution $\boldsymbol{\pi}$ is defined by the Softmax function $\sigma(\cdot)$ and the expected free energy (EFE) of the policies, denoted by $\mathbf{G}$. The EFE incorporates goal-directed components of action selection. Additionally, the precision term $\gamma$ plays a role in encoding the confidence of beliefs concerning $\mathbf{G}$.
@@
$\mathbf{G}\left(\pi_{t}\right)=\sum_{\tau=t+1}^{K}D_{KL}\left[Q\left(o_{\tau}|\pi\right)\parallel P\left(o_{\tau}\right)\right]+\mathbb{E}_{\tilde{Q}}\left[H\left[P\left(o_{\tau}|s_{\tau}\right)\right]\right]$ (4)
@@
After scoring all the policies using the EFE, action selection is performed by drawing over the action posterior expectations derived from the sufficient statistic $\boldsymbol{\pi}$ computed via (3b). Then, the selected action is executed, the model receives a novel observation and the perception-action cycle starts again. See (Parr et al., 2022) for more details.
@@ 1.3 Hierarchical Active Inference model for reading
  • 3.
    the feedback response $r\in O^{(i)}_{3}$, (“correct” or “wrong”) that reports whether the currently inferred content $C^{(i)}$ has been correctly classified (i.e., assigned to the correct topic) or not. We only use this observation at Level 3 (i.e., $O^{3}_{3}$) of Simulation 4, in order to compare models with or without prior information about the topic of the sentence.
@@
  • 3.
    $\mathbf{A}^{(i)}_{3}$ is a 4-order tensor mapping the hidden states $S^{(i)}$ to the feedback response in $O^{(i)}_{3}$. In our simulation we adopt a binary probability in our simulation: the likelihood is set to 1 if the content belongs to the correct topic and to 0 otherwise. Please note that we only use this factor in Simulation 4.
@@
  • 1.
    $\mathbf{B}^{(i)}_{1}$ maps the transition between content states at successive time steps. Note that in our simulations, we use transitions models without noise (for the control model) and with noise added at one or more hierarchical levels (for the dyslexic models). For this, when we formalize the mapping $p(c^{(i)}_{t}|c^{(i)}_{t-1},\pi_{t})=\delta^{(i)}$ we use a modulating factor $\delta^{(i)}$, which we set to 1 in the case without noise and to a number below 1 in the case with noise. Noise implies that the prior knowledge about (or the working memory) of the content variable fades away during reading – requiring more evidence and more saccades to solve the reading tasks.
  • 2.
@@
  • 3.
    $\mathbf{C}_{3}^{(i)}$, prior preferences on the feedback response $O_{3}^{(i)}$ (e.g. for Level 3 encoding sentences, a higher prior for having classified the sentence according to the ”correct” topic than the ”wrong” topic. Please note that we only use this factor in Simulation 4.
@@ 1.4 Saccade selection
Policy selection considers the Expected Free Energy (EFE) described in (4). In our simulations, we employ flat prior preferences on Content and Location (denoted as $\mathbf{C}^{(i)}_{1}$ and $\mathbf{C}^{(i)}_{2})$, respectively), resulting in policy selection that solely weighs the information gain associated with the following word, syllable, or letter. However, in Simulation 4, policy selection also depends on the greater preference $\mathbf{C}^{(i)}_{3})$ for receiving a ”correct” than a ”wrong” feedback response (we remind that this factor is absent in the other simulations).
@@
Furthermore, policy selection considers the prior distribution over policies (denoted as $\mathbf{E}^{(i)}$). Our simulations adopt flat priors for Level 1 ($\mathbf{E}^{(1)}$) and Level 2 ($\mathbf{E}^{(2)}$), signifying no inherent constraints on transitioning between letters within a syllable or between syllables within a word. To discourage overly long saccades, a finite moving window for favored transitions among word locations is set for Level 3 ($\mathbf{E}^{(3)}$). This is achieved through a Poisson distribution $Pois(\lambda)$, where $\lambda$ equals 6 (as displayed in Figure 2, depicted by red bars). Additionally, the probability of reading the same word consecutively is set to zero, resulting in the normalization of the distribution. This culminates in the distribution depicted by the black bars in Figure 2. The chosen distribution enables the creation of a finite window to determine probabilities related to the content under consideration. Iterating this window at each step enables the processing of sentences of arbitrary length.
@@
Figure 2: Priors on the policies that make transitions between word locations at Level 3 of the generative model. We first set the prior as a Poisson distribution with $\lambda=6$ (red bars), centered on the current location (i.e., the value of the maximum of the distribution is in 0). We next set the probability of reading the same word two consecutive times to zero and re-normalize the distribution, leading the the distribution shown by the black bars.
@@
Furthermore, Figure 3 presents a schematic illustration of the textual elements inferred by the model while reading the sentence ”Active inference model of eye movements” as depicted in Figure 1. The diagram distinguishes between two factors of the generative model – Content and Location – illustrated by the left and right trees, respectively. The left (content) tree in Figure 3 demonstrates that the model infers the first word ”active” at Level 2 by examining the letter ”a”. This, in turn, leads to the inference of the syllable ”ac” at Level 1, followed by the letters ”t” and ”e”, enabling the inference of the syllable ”tive” at Level 2. The right (location) tree in Figure 3 shows the respective locations of these three letters ”a”, ”t”, and ”e” within their syllables at Level 1, and the positions of ”ac” and ”tive” within the word at Level 2. This process continues for the subsequent words (excluding ”of” and ”eye”, which are skipped during reading, as seen in Figure 3), until the entire sentence is recognized.
@@
Figure 3: Graphical explanation of the model’s reading process for the example sentence ”Active inference model of eye movements” introduced in Figure 1. The figure illustrates the distinction between two factors of the generative model: content (left tree) and location (right tree). Refer to the main text for a detailed explanation.
@@ 2.1 Simulation 1: Reading single words
In this task, the objective is to read and recognize 100 words consisting of either 4 or 8 letters, selected from a pool of 726 words ranging from 1 to 8 letters, sourced from the BERT dictionary (see Tab. S.14 for the list of words). For ease of evaluation, each word is assigned the same a-priori probability of 1/726 in the model.
@@
Previous research has demonstrated that individuals with dyslexia exhibit a reading style characterized by more fragmented and laborious patterns of eye movement, with a higher number of shorter saccades when reading words, and these difficulties are more pronounced with longer words (Zoccolotti et al., 2005; MacKeben et al., 2004; Franzen et al., 2021). To assess whether our model accurately replicates these findings, we compare two versions: (i) a Control model (CM) with no noise in its transition functions, allowing it to adequately consider the prior context during reading, and (ii) a Dyslexic model (DM), where we introduce noise in the transition functions of both the first (syllable) level and the second (word) levels, in line with the hypothesis that dyslexics struggle to integrate prior linguistic context (Jaffe-Dax et al., 2015).
@@
Moreover, the DM exhibits a significantly higher number of total saccades (Figure 5(a)) and backward saccades (Figure 5(b)) while reading both 4- and 8-letter words compared to the CM. Detailed statistical comparisons are available in Tab. S.2 and S.3. To delve deeper into the impact of noise on eye movement dynamics during word reading, we created two variants of the DM, one with noise in the transition functions at the syllable level (DM - Noise on Level 1) and another with noise at the word level (DM - Noise on Level 2). This analysis reveals that noise at the first (syllable) level has a more significant effect than noise at the second (word) level.
Figure 5: Simulation 1 results: Total number of saccades (a) and number of backward saccades (b) while reading 4-letter and 8-letter words. The figure compares the Control Model (CM) with no noise, the Dyslexic Model (DM) with noise at both hierarchical levels, and two versions of the DM, with noise only at the level of syllables (DM - Noise on Level 1) or words (DM - Noise on Level 2).
@@
Figure 6: Simulation 1 results: Amplitude of forward (a) and backward (b) saccades while reading 4-letter and 8-letter words. The figure compares two models: the Control Model (CM) with no noise, and the Dyslexic Model (DM) with noise at both hierarchical levels.
In summary, our model successfully reproduces a wide range of empirical findings regarding how dyslexic individuals read single words, encompassing accuracy, reaction times, and eye movements. The lower probability assigned to words by the DM compared to the CM, as depicted in Figure 4(a), closely mirrors the results reported by (Zoccolotti et al., 2005, Fig. 1(B)). The accuracy results (Table S.1) are also consistent with the finding that dyslexic individuals fail to correctly read a small percentage of single words. Moreover, our simulations reveal that dyslexic individuals read significantly slower than controls (Zoccolotti et al., 2005; O’Brien et al., 2005, Fig. 1(A)). This slower reading arises from the DM making more saccades than the CM, as illustrated in Figure 5, with reading time being proportional to the number of saccades.
Furthermore, the results displayed in Figure 6(a) closely match the finding that dyslexic individuals make shorter forward saccades compared to controls, particularly evident for longer words (MacKeben et al., 2004, Fig. 4). Lastly, the results presented in Figure 6(b) closely resemble the finding that dyslexic individuals make significantly more and larger backward saccades than controls during reading (Prado et al., 2007, Fig. 5). Collectively, these outcomes demonstrate that our model effectively captures possible aspects of reading skill differences and/or development, providing valuable insights into the underlying mechanisms of reading disorders like dyslexia.
Significantly, our model offers a mechanistic understanding for all these observed findings, including the fragmented reading style of dyslexics. The DM’s shorter forward saccades stem from its poor contextual memory, making it less capable of predicting the next word efficiently. Consequently, it requires more saccades and time to read a text. Additionally, the increased need for backward saccades arises because the DM occasionally has to backtrack in the text to retrieve lost context. Moreover, our model readily explains why dyslexic reading impairments are more pronounced for longer words: noise in the transition function(s) accumulates over time, making reading longer words progressively more challenging. This comprehensive mechanistic explanation underscores the utility of our model in shedding light on the underlying cognitive processes contributing to dyslexia and its impact on reading behavior.

### v3

@@ 2.2 Brief introduction to Active Inference
Active Inference is a framework that models an agent’s action-perception loop by minimizing variational free energy (parr2022active, ). The key objective is to minimize free energy, and to achieve this, the agent possesses a generative model (like the one illustrated in Figure 1), which captures the joint probability of the stochastic variables (hidden states and observations) using the formalism of probabilistic graphical models (bishop2006pattern, ). The agent’s perception involves making inferences about the hidden states based on the observed sensory inputs, while action involves selecting actions that can change the hidden states and optimize the model’s predictions. The generative model for active inference is defined as follows:
$\hskip-10.00002ptP\left(o_{0:K},s_{0:K},u_{1:K},\gamma|\boldsymbol{\Theta}\right)=$
$\hskip-10.00002ptP\left(\gamma|\boldsymbol{\Theta}\right)P\left(\pi|\gamma,\boldsymbol{\Theta}\right)P\left(s_{0}|\boldsymbol{\Theta}\right)\prod_{t=0}^{K}P\left(o_{t}|s_{t},\boldsymbol{\Theta}\right)P\left(s_{t+1}|s_{t},\pi_{t},\boldsymbol{\Theta}\right)$ (1)
@@
[…]
  • The matrix $\mathbf{C}$ encodes an a-priori probability distribution over observations $P\left(o_{\tau}\right)$ and typically encodes the agent’s preferences; this prior distribution parameter is needed for the resolution of Equation (5), with $P\left(o_{\tau}\right)=\mathbf{C}$.
@@
In active inference, the process of perception involves estimating hidden states based on observations and previous hidden states. At the start of the simulation, the model has access to an initial state estimate $s_{0}$ through $\mathbf{D}$ and receives an observation $o_{0}$ that helps refine the estimate by using the likelihood matrix $\mathbf{A}$. Subsequently, for each time step $t=1,\dots,K$, the model infers its current hidden state $s_{t}$ by considering the transitions determined by the control state $u_{t}$, as specified in $\mathbf{B}$. Active inference employs an approximate posterior over (past, present, and future) hidden states and parameters ($s_{0:K},u_{1:K},\gamma$). This approach utilizes a factorized form of variational inference, which corresponds to a framework developed in physics known as mean field theory (Parisi:111935, ). Using a mean field approximation, namely assuming that all variables are independent, the factorized approximated posterior can be expressed as:
@@
where the sufficient statistics are encoded by the expectations $\boldsymbol{\mu}=\left(\mathbf{\tilde{s}}^{\boldsymbol{\pi}},\boldsymbol{\pi},\boldsymbol{\gamma}\right)$, with $\mathbf{\tilde{s}}^{\boldsymbol{\pi}}=\mathbf{\tilde{s}}_{0}^{\boldsymbol{\pi}},\dots,\mathbf{\tilde{s}}_{K}^{\boldsymbol{\pi}}$. Following a variational approach, the approximate posterior $Q(s_{t}|\pi_{t})$ over hidden states under a given policy $\pi_{t}$ at each time step $t$ is optimized by minimizing the variational free energy ($\mathbf{F})$ (see (parr2022active, )) defined as:
$\mathbf{F}_{t}=D_{KL}[Q(s_{t}|\pi_{t})\parallel P(s_{t}|\pi_{t})]-\mathbb{E}_{Q(s_{t}|\pi_{t})}[\ln P(o_{t}|s_{t})]$ (3)
The first term, $D_{\mathrm{KL}}\left[Q(s_{t}|\pi_{t})\,\|\,P(s_{t}|\pi_{t})\right]$, is the Kullback-Leibler (KL) divergence between the approximate posterior and the prior over hidden states. This term quantifies the complexity of the posterior—i.e., how much information (in bits or nats) is required to move from prior beliefs $P(s_{t}|\pi_{t})$ to the posterior $Q(s_{t}|\pi_{t})$. Minimizing this term encourages the posterior belief to have low complexity, i.e., to remain close to prior expectations.
The second term, $\mathbb{E}_{Q(s_{t}|\pi_{t})}\left[\ln P(o_{t}|s_{t})\right]$, corresponds to the expected negative log-likelihood (or inaccuracy) of the observation $o_{t}$ given the inferred hidden states. Minimizing this term encourages the posterior beliefs to be predictive of the actual sensory observations, thus promoting accuracy. In sum, minimizing the variational free energy trades off complexity against accuracy.
The approximate posterior $Q(s_{t}|\pi_{t})$ is optimized when the sufficient statistics are:
@@
Action selection in the active inference framework involves choosing a policy, represented by a sequence of control states $u_{1},u_{2},\dots,u_{t}$, that is expected to minimize free energy most effectively in the future. The policy distribution $\boldsymbol{\pi}$ is defined by the Softmax function $\sigma(\cdot)$. In this equation, a crucial rule is played by the expected free energy (EFE) of the policies, denoted by $\mathbf{G}$. The EFE incorporates goal-directed components of action selection and can be interpreted as the expected variational free energy under future outcomes. Additionally, in this equation the precision term $\gamma$ plays a role in encoding the confidence of beliefs concerning $\mathbf{G}$.
@@
$\hskip-10.00002pt\mathbf{G}\left(\pi_{t}\right)=\sum_{\tau=t+1}^{K}D_{KL}\left[Q\left(o_{\tau}|\pi\right)\parallel P\left(o_{\tau}\right)\right]+\mathbb{E}_{\tilde{Q}}\left[H\left[P\left(o_{\tau}|s_{\tau}\right)\right]\right]$ (5)
@@
After scoring all the policies using the EFE, action selection is performed by drawing over the action posterior expectations derived from the sufficient statistic $\boldsymbol{\pi}$ computed via (4b). Then, the selected action is executed, the model receives a novel observation and the perception-action cycle starts again. We refer the reader to (parr2022active, ) for a complete derivation of the equations presented in this section.
@@ 2.3 Hierarchical Active Inference model for reading
  • the feedback response $r\in O^{(i)}_{3}$, (“correct“ or “wrong”) that reports whether the currently inferred content $C^{(i)}$ has been correctly classified (i.e., assigned to the correct topic) or not. We only use this observation at Level 3 (i.e., $O^{3}_{3}$) of Simulation 4, in order to compare models with or without prior information about the topic of the sentence.
@@
  • $\mathbf{A}^{(i)}_{3}$ is a 4-order tensor mapping the hidden states $S^{(i)}$ to the feedback response in $O^{(i)}_{3}$. In our simulation we adopt a binary probability in our simulation: the likelihood is set to 1 if the content belongs to the correct topic and to 0 otherwise. Please note that in this paper this factor is used only in Simulation 4.
@@
  • $\mathbf{B}^{(i)}_{1}$ maps the transition between content states at successive time steps. Note that in our simulations, we use transition models without noise (for the control model) and with noise added at one or more hierarchical levels (for the dyslexic models). For this, we formalize the mapping $p(c^{(i)}_{t}|c^{(i)}_{t-1},\pi_{t})$ by assigning $1-\delta$ to the correct target transition and assigning the value $\delta/(M^{(i)}-1)$ across the other $M^{(i)}-1$ states, where $M^{(i)}$ is the total number of Content states $S_{1}^{(i)}$. This modulating factor $\delta^{(i)}\in[0,1]$, is set to 0 in the control model and to 0.15 in the dyslexic models.
@@
  • $\mathbf{C}_{3}^{(i)}$, prior preferences on the feedback response $O_{3}^{(i)}$ (e.g. for Level 3 encoding sentences, a higher prior for having classified the sentence according to the “correct” topic than the “wrong” topic. Please note that we only use this factor in Simulation 4.
@@ 2.4 Saccade selection
Policy selection considers the Expected Free Energy (EFE) described in (5). In our simulations, we employ flat prior preferences on Content and Location (denoted as $\mathbf{C}^{(i)}_{1}$ and $\mathbf{C}^{(i)}_{2})$, respectively), resulting in policy selection that solely weighs the information gain associated with the following word, syllable, or letter. However, in Simulation 4, policy selection also depends on the greater preference $(\mathbf{C}^{(i)}_{3})$ for receiving a “correct” than a “wrong” feedback response (we remind that this factor is absent in the other simulations).
@@
Furthermore, policy selection considers the prior distribution over policies (denoted as $\mathbf{E}^{(i)}$). Our simulations adopt flat priors for Level 1 ($\mathbf{E}^{(1)}$) and Level 2 ($\mathbf{E}^{(2)}$), signifying no inherent constraints on transitioning between letters within a syllable or between syllables within a word. To discourage overly long saccades, a finite moving window for favored transitions among word locations is set for Level 3 ($\mathbf{E}^{(3)}$). This is achieved through a Poisson distribution $Pois(\lambda)$, where $\lambda$ equals 6 (as displayed in Figure 2, depicted by red bars). Additionally, the probability of reading the same word after having already recognized it is set to zero, resulting in the normalization of the distribution (note however that it the distribution allows backtracking to a previous word during reading). This culminates in the distribution depicted by the black bars in Figure 2. The chosen distribution enables the creation of a finite window to determine probabilities related to the content under consideration. Iterating this window at each step enables the processing of sentences of arbitrary length.
@@
Figure 2: Priors on the policies that make transitions between word locations at Level 3 of the generative model. We first set the prior as a Poisson distribution with $\lambda=6$ (red bars), centered on the current location (i.e., the value of the maximum of the distribution is in 0). We next set the probability of reading the same word two consecutive times to zero and re-normalize the distribution, leading the distribution shown by the black bars.
@@
Furthermore, Figure 3 presents a schematic illustration of the textual elements inferred by the model while reading the sentence “Active inference model of eye movements” as depicted in Figure 1. The diagram distinguishes between two factors of the generative model – Content and Location – illustrated by the left and right trees, respectively. The left (content) tree in Figure 3 demonstrates that the model infers the first word “active” at Level 2 by examining the letter “a”. This, in turn, leads to the inference of the syllable “ac” at Level 1, followed by the letters “t” and “e”, enabling the inference of the syllable “tive” at Level 2. The right (location) tree in Figure 3 shows the respective locations of these three letters “a”, “t”, and “e” within their syllables at Level 1, and the positions of “ac” and “tive” within the word at Level 2. This process continues for the subsequent words (excluding “of” and “eye”, which are skipped during reading, as seen in Figure 3), until the entire sentence is recognized.
@@
Figure 3: Graphical explanation of the model’s reading process for the example sentence “Active inference model of eye movements” introduced in Figure 1. The figure illustrates the distinction between two factors of the generative model: content (left tree) and location (right tree). Refer to the main text for a detailed explanation.
@@ 3.1 Simulation 1: Reading single words
In this task, the objective is to read and recognize 100 words consisting of either 4 or 8 letters, selected from a pool of 726 words ranging from 1 to 8 letters (see Tab. S.14 for the list of words used in this simulation, all of which are part of the dictionary of BERT). Syllables for each word are generated from words using the methods described in (liang1983word, ). For ease of evaluation, each word is assigned the same a-priori probability of 1/726 in the model.
@@
To better evaluate the model, we compare two versions: (i) a Control model (CM) with no noise in its transition functions, allowing it to adequately consider the prior context during reading, and (ii) a Dyslexic model (DM), aiming to capture qualitatively key aspects of this disorder. Previous research has demonstrated that individuals with dyslexia exhibit a reading style characterized by more fragmented and laborious patterns of eye movement, with a higher number of shorter saccades when reading words, and these difficulties are more pronounced with longer words (zoccolotti2005word, ; mackeben2004eye, ; franzen2021individuals, ). To provide a proof of concept that our DM model can replicate this pattern of results, here we follow the proposal that dyslexics struggle to integrate prior linguistic context (jaffe2015computational, ) – and introduce noise in the transition functions of both the first (syllable) level and the second (word) levels. For this, we set $\delta^{(1)}=\delta^{(2)}=0.15$ in the DM model, see Methods section.
@@
Moreover, the DM exhibits a significantly higher number of total saccades (Figure 5(a)) and backward saccades (i.e., saccades to any of the preceding letters, Figure 5(b)) while reading both 4- and 8-letter words compared to the CM. Detailed statistical comparisons are available in Tab. S.2 and S.3. To delve deeper into the impact of noise on eye movement dynamics during word reading, we created two variants of the DM, one with noise in the transition functions at the syllable level (DM - Noise on Level 1) and another with noise at the word level (DM - Noise on Level 2). This analysis reveals that noise at the first (syllable) level has a more significant effect than noise at the second (word) level.
Figure 5: Simulation 1 results: Total number of saccades (a) and number of backward saccades (b) while reading 4-letter and 8-letter words. The figure compares the Control Model (CM) with no noise, the Dyslexic Model (DM) with noise at all hierarchical levels, and two versions of the DM, with noise only at the level of syllables (DM - Noise on Level 1) or words (DM - Noise on Level 2). See Table S.2 and Table S.3 for detailed results and statistical comparisons.Horizontal bars indicate statistically significant differences from the Control Model (CM). Significance levels are indicated as follows: * $p<0.05$, ** $p<0.01$, *** $p<0.001$.
@@
Figure 6: Simulation 1 results: Amplitude of forward (a) and backward (b) saccades while reading 4-letter and 8-letter words. The figure compares two models: the Control Model (CM) with no noise, and the Dyslexic Model (DM) with noise at both hierarchical levels. See Table S.4 and Table S.5 for detailed results and statistical comparisons.
In summary, our model successfully reproduces a wide range of empirical findings regarding how dyslexic individuals read single words, encompassing accuracy, reaction times, and eye movements. The lower probability assigned to words by the DM compared to the CM, as depicted in Figure 4(a), closely mirrors the results reported by (zoccolotti2005word, , Fig. 1(B)). The accuracy results (Table S.1) are also consistent with the finding that dyslexic individuals fail to correctly read a small percentage of single words. Moreover, our simulations reveal that dyslexic individuals read significantly slower than controls (zoccolotti2005word, ; o2005effect, , Fig. 1(A)). This slower reading arises from the DM making more saccades than the CM, as illustrated in Figure 5, with reading time being proportional to the number of saccades.
Furthermore, the results displayed in Figure 6(a) closely match the finding that dyslexic individuals make shorter forward saccades compared to controls, particularly evident for longer words (mackeben2004eye, , Fig. 4). Lastly, the results presented in Figure 6(b) closely resemble the finding that dyslexic individuals make significantly more and larger backward saccades than controls during reading (prado2007eye, , Fig. 5). Collectively, these outcomes demonstrate that our model effectively captures possible aspects of reading skill differences and/or development, providing valuable insights into the underlying mechanisms of reading disorders like dyslexia.
Significantly, our model offers a mechanistic understanding for all these observed findings, including the fragmented reading style of dyslexics. The DM’s shorter forward saccades stem from its poor contextual memory, making it less capable of predicting the next word efficiently. Consequently, it requires more saccades and time to read a text. Additionally, the increased need for backward saccades arises because the DM occasionally has to backtrack in the text to retrieve lost context. Moreover, our model readily explains why dyslexic reading impairments are more pronounced for longer words: noise in the transition function(s) accumulates over time, making reading longer words progressively more challenging. This comprehensive mechanistic explanation underscores the potential usefulness of our model in shedding light on the underlying cognitive processes contributing to dyslexia and its impact on reading behavior.